If plywood comes in pieces that measure 8 feet by 4 feet, how many pieces of plywood are needed to cover the roof? Assume tht both halves of the roof are the same.
Roof is 42ft by 25ft

Answer by richwmiller(17219)


In theory 42*25=1050 sq feet per side
8*4=32 sq ft per sheet
1050/32=32.8125 sheets per side but it would involve cutting to fit.
for both sides of the roof
65.625 sheets total